The following Payroll Journal entries for Oct. 15 were made by your predecessor. For FICA tax, assume that the social security rate is 6.0% and the Medicare rate is 1.5%. The state and federal unemployment tax rates are 5.4% and 0.8%, respectively. The company offers 401k plans to employees. Review the journal entries, then answer the questions that follow.

PAGE 32

JOURNAL

ACCOUNTING EQUATION

DATE DESCRIPTION POST. REF. DEBIT CREDIT ASSETS LIABILITIES EQUITY

1

Oct. 15

Sales Salaries Expense

93,000.00

2

Officers Salaries Expense

334,800.00

3

Office Salaries Expense

62,000.00

4

Factory Wages Expense

130,200.00

5

Social Security Tax Payable

37,200.00

6

Medicare Tax Payable

9,300.00

7

Employees Federal Income Tax Payable

111,600.00

8

Medical Insurance Payable

68,200.00

9

Retirement Contributions Payable

93,000.00

10

Salaries Payable

300,700.00

11

15

Payroll Tax Expense

47,833.00

12

Social Security Tax Payable

37,200.00

13

Medicare Tax Payable

9,300.00

14

Federal Unemployment Tax Payable

172.00

15

State Unemployment Tax Payable

1,161.00

16

15

Pension Expense

31,000.00

17

Cash

31,000.00

In order to confirm the previous clerk’s payroll calculations, you have been asked to supply the following amounts based on your review of the payroll entries. These amounts will be checked against the company records and investigated further if necessary.

1. Determine the payroll amount subject to federal and state unemployment taxes in this payroll.

2. What is the total payroll for Copperfield and Company shown in these journal entries?

3. What is Copperfield and Company’s share of FICA taxes in this payroll?

4. How much has Copperfield and Company contributed to employee 401k plans in this payroll?
Answer #1

PART – 1)

Payroll subject to the federal and state unemployment taxes = Federal unemployment tax/Tax rates

= (172/0.8%)

= 21,500

Or

= State unemployment tax payable/Tax rate

= 1,161/5.4%

= 21,500                                                                           

PART – 2)

Total payroll = Officer salary + Sales salary + Office salary + Factory wages

= 334,800 + 93,000 + 62,000 + 130,200

= 620,000

PART – 3)

FICA taxes in payroll = Medical tax payable + Social security tax payable

                                    = 9,300 + 37,200
                                    = 46,500

PART – 4)

Contribution to employee’s 401k plan = Pension payable

= 31,000

The contribution to retirement and medical insurance is made by employees.